# Uncertainty Estimation and Quantification for LLMs: A Simple Supervised Approach

### v1

@@ 5 Conclusions
In this paper, we study the problem of uncertainty estimation and calibration for LLMs. We follow a simple and standard supervised idea and use the labeled NLP datasets to train an uncertainty estimation model for LLMs. Our finding is that, first, the proposed supervised methods have better performances than the existing unsupervised methods. Second, the hidden activations of the LLMs contain uncertainty information about the LLMs’ responses. Third, the black-box regime of our approach (Bb-S) provides a new approach to estimating the uncertainty of closed-source LLMs. Lastly, we distinguish the task of uncertainty estimation from uncertainty calibration and show that a better uncertainty estimation model leads to better calibration performance. We also remark on the following two aspects:
One limitation of our proposed supervised method is that it critically relies on the labeled data. For the scope of our paper, we restrict the discussion to the NLP tasks and datasets. One future direction is to utilize the human-annotated data for LLMs’ responses to train a supervised uncertainty estimation model for open-question prompts. We believe the findings that the supervised method gives a better performance and the hidden activations contain the uncertainty information will persist.
Appendix A Details for the Numerical Experiments
A.1 Details for used features
In Section 3.1, we briefly discussed how to construct features $\bm{v}_{i}$ for the uncertainty prediction based on the original prompt response pair $(\bm{x}_{i},\bm{y}_{i}).$ In Table 5, we provide a full list of the probability- and entropy-related features used in our uncertainty estimation model where $p_{\theta}$ denotes the LLM as previous.
@@ A.3 Details of the training procedure
In the experiment section of the main text, the features in the Wb-S for natural language generation tasks include (i) all the features used in the Gb-S, (ii) the hidden activations of the last token of the question from the middle layer (LLaMA2-7B: 16th layer; Gemma-7B: 14th layer), and (iii) the hidden activations of the last token of the answer from the middle layer. Therefore, in these natural language generation tasks, the dimension is 8212 for LLaMA2-7B and 6164 for Gemma-7B.
The features in the Wb-S for the multiple-choice task include (i) all the features used in the Gb-S and (ii) the hidden activations of the last token of the answer (letter A, B, C, or D) from the middle layer. The dimension is 4101 for LLaMA2-7B and 3077 for Gemma-7B.
Notably, there are many choices of the hidden activations employed in the Wb-S. Besides what has been shown in Section 4.3, we provide further discussion in Section B.
Bb-S. The idea of building a supervised calibration model for a black-box LLM is to use the hidden layers and output distributions from another open-source LLM model by feeding it with the question and the provided response. Therefore, the features available for the Wb-S are also available for the open-source LLM, so we just take the corresponding features from the open-source LLM in the Bb-S. Hence, in the natural language generation tasks, the input dimension of the calibration model is 4196 (including hidden activations of the question and answer and 20 entropy and likelihood-related features, $2\times 2048+20$) for Gemma-2B, 6164 for Gemma-7B, 8212 for LLaMA2-7B, and 10260 for LLaMA2-13B. In the multiple-choice task, the dimension is 2053 for Gemma-2B (including the hidden activations of the answer and 5 entropy- and probability-related features used in the Gb-S), 3077 for Gemma-7B, 4101 for LLaMA2-7B, and 5125 for LLaMA2-13B.
Appendix B Additional results and visualizations
In Section 4.3, we show the advantage of utilizing the hidden activations of the answer from the middle layer of the LLM to estimate the uncertainty in Wb-S. In this section, we further discuss the impact of employing the hidden activations from the question in the Wb-S.
@@ Appendix B Additional results and visualizations
Question’s hidden activation cannot help to generate more uncertainty information Again from Figure 5, by comparing the performance of Answer and Question-Answer, we find that the inclusion of question’s activation has little impact on improving the performance. This shows that the uncertainty from the question might have already been well encoded in the last token activation of the answer. In other words, it justifies our approach to employ Wb-S, which includes the answer’s hidden activation and the entropy- and probability-related features.
(a) Features from Gemma-7B
(b) Features from LLaMA2-7B
Figure 5: Performance comparison of using last-token middle layer hidden activations of the answer (Answer) or the concatenation of the question and answer (Question-Answer) as features in the Wb-S, where the features in Gb-S are also included in Wb-S. In the natural language generation tasks, the dimensions of Gb-S, Question, Answer, and Question-Answer for Gemma-7B are 20, 3082, 3092, and 6164, while for LLaMA2-7B they are 20, 4106, 4116, and 8212, respectively. In the MMLU task, for Gemma-7B they are 5, 3077, 3077, and 6149, while for LLaMA2-7B they are 5, 4101, 4101, and 8197, respectively.
Figure 6: More distribution of values from specific neurons on TriviaQA dataset from Gemma-7B when the model is generating answers. The plots are obtained in the same way as Figure 4. This again proves the validity of incorporating hidden activations from answers.
The middle layer is still better than the last layer. In Section 4.3, Figure 1 shows that when using the hidden activation of the answer in the Wb-S, the middle layer of the LLM is a better choice than the last layer. The next question is: Does this conclusion still hold for using the concatenated hidden activations of the question and answer? We depict the experiment result in Figure 8, which is consistent with the conclusion drawn from Figure 1.
@@
(a) Features from Gemma-7B
(b) Features from LLaMA2-7B
Figure 8: Performance comparison of using question-answer concatenated hidden activations from different tokens and layers as features in the Wb-S method. Scores are normalized in [0,1], where a lower value indicates larger uncertainty. For Gemma-7B, the dimension of the Wb-S input is 6164 (3072 from the question, 3072 from the answer, and 20 from the grey-box features). For LLaMA2-7B, it is 8212.
Appendix C Examples

### v2

@@ 5 Conclusions
In this paper, we study the problem of uncertainty estimation and calibration for LLMs. We follow a simple and standard supervised idea and use the labeled NLP datasets to train an uncertainty estimation model for LLMs. Our finding is that, first, the proposed supervised methods have better performances than the existing unsupervised methods. Second, the hidden activations of the LLMs contain uncertainty information about the LLMs’ responses. Third, the black-box regime of our approach (Bb-S) provides a new approach to estimating the uncertainty of closed-source LLMs. Lastly, we distinguish the task of uncertainty estimation from uncertainty calibration and show that a better uncertainty estimation model leads to better calibration performance. One limitation of our proposed supervised method is that it critically relies on the labeled data. For the scope of our paper, we restrict the discussion to the NLP tasks and datasets. One future direction is to utilize the human-annotated data for LLMs’ responses to train a supervised uncertainty estimation model for open-question prompts. We believe the findings that the supervised method gives a better performance and the hidden activations contain the uncertainty information will persist.
Appendix A Interpreting the Uncertainty Estimation
A.1 Layer comparison
For general LLMs, each token is associated with a relatively large number of hidden layers (32 layers for LLaMA2-7B for example), each of which is represented by high-dimensional vectors (4096 for LLaMA2-7B). Thus it is generally not a good practice to incorporate all hidden layers as features for the uncertainty estimation due to this dimensionality. Previous works find that the middle layer and the last layer activations of the LLM’s last token contain the most useful features for supervised learning (Burns et al., 2022; Chen et al., 2024; Ahdritz et al., 2024; Azaria and Mitchell, 2023). To investigate the layer-wise effect for uncertainty estimation, we implement our Wb-S method with features different in two aspects: (i) different layers within the LLM architecture, specifically focusing on the middle and last layers (e.g., LLaMA2-7B and LLaMA3-8B: 16th and 32nd layers out of 32 layers with 4096 dimensions; Gemma-7B: 14th and 28th layers out of 28 layers with 3072 dimensions); and (ii) position of token activations, including averaging hidden activations over all the prompt/answer tokens or utilizing the hidden activation of the last token. The second aspect makes sense when the output contains more than one token, so we conduct this experiment on the natural language generation tasks only. Figure 1 gives a visualization of the comparison result. While the performances of these different feature extraction ways are quite similar in terms of performance across different tasks and LLMs, activation features from the middle layer generally perform better than the last layer. This may come from the fact that the last layer focuses more on the generation of the next token instead of summarizing information of the whole sentence, as has been discussed by Azaria and Mitchell (2023).
A.2 Scaling effect
Figure 2: (Left) Using the hidden activations of LLaMA2-7B and LLaMA2-13B to estimate the uncertainty of the answer provided by Gemma-7B. (Middle) Using the hidden activations of Gemma-2B and Gemma-7B to estimate the uncertainty of the answer provided by LLaMA2-7B. (Right) Using the hidden activations of Gemma-2B and Gemma-7B to estimate the uncertainty of the answer provided by LLaMA3-8B
We find that larger LLM does encode better knowledge about the uncertainty, which is attributed to their improved knowledge in answering the questions. We also note that in the case of using Gemma to predict LLaMA2-7B, even a small tool LLM (Gemma-2B) is capable of achieving better performance than the Gb-S that only uses the entropy- and probability-related features from the target LLM. This result also underscores the benefits of adopting the internal state in estimating the uncertainty, even from an LLM different from the one generating the answers.
A.3 Histogram of correlations
Figure 3 plots the histograms of the pairwise correlations between the neuron activations and the labels (whether the LLM’s response is correct). We make two observations here: First, for both LLMs, some neurons have a significantly positive (or negative) correlation with the label. We can interpret these neurons as the uncertainty neuron for the corresponding task. When these neurons are activated, the LLMs are uncertain about their responses. Second, Gemma-7B and LLaMA3-8B have more significant neurons than LLaMA2-7B, and this is consistent with the better performance of Gemma-7B and LLaMA3-8B in Table 1 and Table 2. Also, this reinforces that the hidden activations of the LLMs contain uncertainty information about the LLM’s output.
Figure 3: The histograms of the pairwise correlations on the TriviaQA task between the neuron activations and the labels (whether the LLM’s response is correct), where the neural values are the last-token hidden activations of answers from the middle layer (upper) and the last layer (lower) of two models respectively.
Figure 4 plots some example neurons’ activation by selecting the neurons with the largest absolute correlations in Figure 3. More neurons from the last layer can be found in Figure 5. These neurons as an individual indicator exhibit different distributional patterns when the response is correct compared to when the response is incorrect, and thus reflect the uncertainty of the LLM’s responses.
Figure 4: Distribution of values from particular neurons of mid-layers on TriviaQA dataset.
Figure 5: More distribution of values from specific neurons of last layers on the TriviaQA dataset. The plots are obtained in the same way as Figure 4.
A.4 Proof of Proposition 3.1
The proof of Proposition 3.1 follows from the definition of $f^{*}.$
Appendix B Calibration performance
In Section 3.2, we distinguish the two tasks of uncertainty estimation and uncertainty calibration. Throughout the paper, we have been focused on improving the performance on the task of uncertainty estimation – to predict when the LLM is uncertain about its response. Generally, a better uncertainty estimation model leads to one with better calibration performance. The calibration (or recalibration) of the uncertainty estimation model can be indeed reduced to the classic ML setting which does not involve the LLM. Table 4 gives the calibration performance and we see an advantage of our supervised methods over benchmark methods consistent with the AUROC performance in Table 1. We adopt the histogram binning method here because we find that the temperature scaling method and the Platt scaling method will give all predicted scores concentrated within a small range such as $[0.2,0.6]$. We also do not exclude the possibility that the other calibration methods can give even better performance. The point to make here is that uncertainty estimation and uncertainty calibration are two closely related tasks. Note that (i) a better uncertainty estimation model leads to a better calibration performance and (ii) the LLMs are pretrained and not designed for these NLP tasks in the first place (see Section 3.3) so that there is no uncertainty score readily available (as the predicted probabilities for the image classifiers); we emphasize the importance of an extra uncertainty estimation procedure as our supervised one so to extract the uncertainty information from the inside of the LLMs.
Metric Dataset Model Benchmarks Ours
MaxL AvgL MaxE AvgE SU A4C Bb-S Gb-S Wb-S
NLL TriviaQA G-7B 0.478 0.500 0.428 0.472 0.739 8.710 0.414 0.467 0.392
L-7B 1.155 0.551 0.575 0.600 1.481 21.119 0.338 0.580 0.388
L-8B 0.483 0.407 0.383 0.401 0.719 8.515 0.423 0.467 0.365
CoQA G-7B 0.778 0.474 0.469 0.476 0.632 8.106 0.474 0.497 0.457
L-7B 1.047 0.620 0.637 0.649 1.358 11.708 0.417 0.607 0.457
L-8B 0.823 0.502 0.508 0.499 0.762 8.007 0.551 0.535 0.507
WMT-14 G-7B 9.674 1.266 0.809 0.618 0.701 17.933 0.454 0.463 0.449
L-7B 1.204 1.150 0.718 0.809 0.796 16.913 0.553 0.622 0.583
L-8B 1.490 0.752 0.652 0.676 0.722 21.340 0.649 0.673 0.612
ECE TriviaQA G-7B 0.152 0.138 0.066 0.115 0.275 0.253 0.056 0.075 0.067
L-7B 0.437 0.068 0.048 0.146 0.188 0.616 0.043 0.087 0.049
L-8B 0.171 0.082 0.046 0.081 0.196 0.283 0.107 0.087 0.075
CoQA G-7B 0.356 0.054 0.112 0.064 0.221 0.237 0.121 0.129 0.113
L-7B 0.397 0.065 0.105 0.073 0.174 0.494 0.052 0.071 0.038
L-8B 0.339 0.031 0.071 0.033 0.196 0.312 0.156 0.110 0.122
WMT-14 G-7B 0.499 0.464 0.234 0.197 0.072 0.521 0.097 0.063 0.073
L-7B 0.164 0.389 0.065 0.269 0.127 0.491 0.045 0.090 0.101
L-8B 0.318 0.192 0.051 0.142 0.029 0.618 0.145 0.201 0.137
Brier TriviaQA G-7B 0.282 0.221 0.224 0.215 0.344 0.279 0.266 0.288 0.282
L-7B 0.431 0.241 0.271 0.259 0.322 0.645 0.334 0.322 0.315
L-8B 0.262 0.192 0.204 0.188 0.291 0.373 0.258 0.265 0.255
CoQA G-7B 0.318 0.174 0.188 0.171 0.232 0.241 0.207 0.218 0.212
L-7B 0.395 0.233 0.242 0.230 0.265 0.464 0.296 0.256 0.276
L-8B 0.338 0.197 0.201 0.191 0.255 0.359 0.258 0.242 0.248
WMT-14 G-7B 0.505 0.454 0.330 0.319 0.247 0.606 0.327 0.287 0.309
L-7B 0.313 0.413 0.271 0.334 0.275 0.502 0.296 0.277 0.288
L-8B 0.343 0.279 0.250 0.263 0.246 0.620 0.282 0.300 0.284
Appendix C Details for the Numerical Experiments
We ran all of our experiments on an AMD EPYC 7452 128-core processor with 4$\times$48G NVIDIA A6000 GPUs.
C.1 Dataset preparation
In the following we provide more information for the three tasks considered in our numerical experiments.
[…]
C.2 Details of the training procedure
@@
In the experiment section of the main text, the features in the Wb-S for natural language generation tasks include (i) all the features used in the Gb-S, (ii) the hidden activations of the last token of the question from the middle layer (LLaMA2-7B or LLaMA3-8B: 16th layer; Gemma-7B: 14th layer), and (iii) the hidden activations of the last token of the answer from the middle layer. Therefore, in these natural language generation tasks, the dimension is 8212 for LLaMA2-7B/LLaMA3-8B and 6164 for Gemma-7B.
The features in the Wb-S for the multiple-choice task include (i) all the features used in the Gb-S and (ii) the hidden activations of the last token of the answer (letter A, B, C, or D) from the middle layer. The dimension is 4101 for LLaMA2-7B/LLaMA3-8B and 3077 for Gemma-7B.
Notably, there are many choices of the hidden activations employed in the Wb-S. Besides what has been shown in Section A, we provide further discussion in Section D.
Bb-S. The idea of building a supervised calibration model for a black-box LLM is to use the hidden layers and output distributions from another open-source LLM model by feeding it with the question and the provided response. Therefore, the features available for the Wb-S are also available for the open-source LLM, so we just take the corresponding features from the open-source LLM in the Bb-S. Hence, in the natural language generation tasks, the input dimension of the calibration model is 4196 (including hidden activations of the question and answer and 20 entropy and likelihood-related features, $2\times 2048+20$) for Gemma-2B, 6164 for Gemma-7B, 8212 for LLaMA2-7B/LLaMA3-8B, and 10260 for LLaMA2-13B. In the multiple-choice task, the dimension is 2053 for Gemma-2B (including the hidden activations of the answer and 5 entropy- and probability-related features used in the Gb-S), 3077 for Gemma-7B, 4101 for LLaMA2-7B/LLaMA3-8B, and 5125 for LLaMA2-13B.
For all these methods, we employ the random forest (Breiman, 2001) using the implementation from the scikit-learn package (Pedregosa et al., 2011) to estimate the uncertainty. The hyperparameters are set as [n_estimators=150, random_state=0, max_depth=8, verbose=2, max_features=45] if the number of selected features is no less than 100 and [n_estimators=100, random_state=0, max_depth=4, verbose=2] otherwise.
Appendix D Additional results and visualizations
In Section A, we show the advantage of utilizing the hidden activations of the answer from the middle layer of the LLM to estimate the uncertainty in Wb-S. In this section, we further discuss the impact of employing the hidden activations from the question in the Wb-S.
@@ Appendix D Additional results and visualizations
Question’s hidden activation cannot help to generate more uncertainty information Again from Figure 6, by comparing the performance of Answer and Question-Answer, we find that the inclusion of question’s activation has little impact on improving the performance. This shows that the uncertainty from the question might have already been well encoded in the last token activation of the answer.
Figure 6: Performance comparison of using last-token middle layer hidden activations of the answer (Answer) or the concatenation of the question and answer (Question-Answer) as features in the Wb-S, where the features in Gb-S are also included in Wb-S. In the natural language generation tasks, the dimensions of Gb-S, Question, Answer, and Question-Answer for Gemma-7B are 20, 3082, 3092, and 6164, while for LLaMA2-7B or LLaMA3-8B they are 20, 4106, 4116, and 8212, respectively. In the MMLU task, for Gemma-7B they are 5, 3077, 3077, and 6149, while for LLaMA2-7B or LLaMA3-8B, they are 5, 4101, 4101, and 8197, respectively.
The middle layer is still better than the last layer. In Section A, Figure 1 shows that when using the hidden activation of the answer in the Wb-S, the middle layer of the LLM is a better choice than the last layer. The next question is: Does this conclusion still hold for using the concatenated hidden activations of the question and answer? We depict the experiment result in Figure 7, which is consistent with the conclusion drawn from Figure 1.
Figure 7: Performance comparison of using question-answer concatenated hidden activations from different tokens and layers as features in the Wb-S method. Scores are normalized in [0,1], where a lower value indicates larger uncertainty. For Gemma-7B, the dimension of the Wb-S input is 6164 (3072 from the question, 3072 from the answer, and 20 from the grey-box features). For LLaMA2-7B/LLaMA3-8B, it is 8212.
@@
Appendix E Examples